Use the information below to answer questions 3.8 to 3.10.

If 𝐶̅= R25 billion, I = R35 billion and c = 0.80 

3.8 The marginal propensity to save is …

[1] 0.80

[2] 1

[3] 1.20

[4] 0.20

3.9 The value of the multiplier is …

[1] 5

[2] 0.20

[3] 1

[4] 2

3.10 The equilibrium level of income is …

[1] 60

[2] 120

[3] 300

[4] 90


1
Expert's answer
2021-07-19T10:33:42-0400

3.8 The marginal propensity to save is:

"mps = 1 - mpc = 1 - 0.8 = 0.2."

So, the correct answer is [4].

3.9 The value of the multiplier is: "m = 1\/mps = 1\/0.2 = 5."

So, the correct answer is [1].

3.10 The equilibrium level of income is:

Y = C + I = 25 + 35 = 60.

So, the correct answer is [1].